OpenStudy (anonymous):

A company is considering implementing one of two quality control plans for monitoring the weights of automobile batteries that it manufactures. If the manufacturing process is working properly, the battery weights are approximately normally distributed with a specified mean and standard deviation. Quality control plan A calls for rejecting a battery as defective if its weight falls more than 2 standard deviations below the specified mean. Quality control plan B calls for rejecting a battery as defective if its weight falls more than 1.5 interquartile ranges below the lower quartile of the specified population. Assume the manufacturing process is under control. a) What proportion of batteries will be rejected by plan A? b) What is the probability that at least 1 of 2 randomly selected batteries will be rejected by plan A? c) What proportion of batteries will be rejected by plan B?

OpenStudy (anonymous):

@jigglypuff314 @Michele_Laino @nincompoop @robtobey

OpenStudy (michele_laino):

for questions a) and c), we have to use the tables about the gaussian distribution

OpenStudy (michele_laino):

have you got those tables?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

now what

OpenStudy (michele_laino):

Now, in order to answer to question a), you have to look at the row, with z=-2.0 and pick the number at the first column

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

.0228

OpenStudy (michele_laino):

ok! Then, the requested probability is: 2.28%

OpenStudy (anonymous):

ok! then what do we do

OpenStudy (michele_laino):

Now, in order to answer to question b), we have to lokk at the row with z= -1.5, and then we have to pick the number at the first column

OpenStudy (michele_laino):

oopss. question c)

OpenStudy (anonymous):

wait so thats the answer for a?

OpenStudy (michele_laino):

no, since we have a different value for z

OpenStudy (anonymous):

wait so we haven't figured out a yet?

OpenStudy (michele_laino):

I think that the requested proportion, is: \[\large \frac{{2.28}}{{100}} \cong \frac{2}{{100}}\]

OpenStudy (anonymous):

im confused

OpenStudy (michele_laino):

using the plan A, we reject 2 batteries every 100 bateries produced

OpenStudy (michele_laino):

oops.. batteries*

OpenStudy (anonymous):

oh okay so 2 out of every 100 batteries are rejected for A

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay :)

OpenStudy (michele_laino):

now, for questio0n c), we have to lookk at row with z=-1.5, and we have to pick the number at the first column, similarly for question a)

OpenStudy (michele_laino):

look*

OpenStudy (anonymous):

0.0668

OpenStudy (michele_laino):

ok! So we can say that since we have: \[\large \frac{{6.68}}{{100}} \cong \frac{7}{{100}}\] then 7 batteries out for every 100 batteries are rejected for plan B

OpenStudy (anonymous):

so for C we say for every 100 batteries for plan B 7 batteries are rejected

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay now b? :)

OpenStudy (michele_laino):

please wait I'm trying to figure out that question...

OpenStudy (anonymous):

thats fine!

OpenStudy (michele_laino):

Sincerely speaking, I don't know that answer

OpenStudy (michele_laino):

Nevertheless, I will continue to try to solve that question

OpenStudy (anonymous):

no its fine i think i got it! could you confirm a different question instead?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

this is the question A GE light bulb is supposed to last for 1200 hours. In fact, light bulbs of this type last only 1185 hours with a standard deviation of 70 hours. What is the probability that a sample of 100 light bulbs will have an average life of at least 1200 hours?

OpenStudy (anonymous):

this my work/ answer 70/sqrt100 = 7 P(xbar>1200)=P(xbar>1200)=(Pz>(1200-1185)/7) = 2.142 normalcdf(1200,1E99,1185,7) = -.9839377721 1+ -.9839377721 = 0.160622279

OpenStudy (michele_laino):

yes! your work is correct, since you have weel apllied the definition of the meaning of the gaussian curve

OpenStudy (michele_laino):

oops..well*

OpenStudy (michele_laino):

please wait, the standard deviation is 70, not 7

OpenStudy (anonymous):

but don't you do 70/sqt100 = 7?

OpenStudy (michele_laino):

you are right, yes we have to keep the standard deviation of the mean

OpenStudy (michele_laino):

ok! your answer is right! more explanation: the probability that a light bulb will have an average life of at least of 1220 hours is: 50-48.34= 1.66%

OpenStudy (michele_laino):

oops..1200 hours

OpenStudy (anonymous):

if i submit just what i said is that good or should i add your further explanation

OpenStudy (michele_laino):

I think that would be better if you express your result as a percentage, since your questions requests a probability. Generally, a probability is expressed in percentage form

OpenStudy (anonymous):

so approx 16%

OpenStudy (michele_laino):

please, wait, I got this: the probability is: 50-48.38=1.62 approximately

OpenStudy (michele_laino):

namely 1.62%

OpenStudy (anonymous):

where did you get the 50 and 48.32

OpenStudy (michele_laino):